OpenStudy (anonymous):

solve the equation. ((4)/(7 x))+((1)/(2x)) = -((1)/(4)) I'm stumped

OpenStudy (anonymous):

\[\frac{4}{7x}+\frac{1}{2x} = -\frac{1}{4}\] First keep in mind that anything times 1 equals itself, e.g. x*1=x Also keep in mind that anything divided by itself also equals 1, e.g. 1 = (5/5) What you want to do is to multiply 1 (in one form or another) by each fraction so that the denominator is the same for all three fractions.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problemo